20% of items produced from a factory are defective.find the probability that in a sample of 5cjosen at random 1; none is defective 2; one is defective 3: p(1
Answer by dkppathak(439)

20% of items produced from a factory are defective.find the probability that in a sample of chosen at random 1; none is defective 2; one is defective 3: p(1 solution
P=0.2 q=0.8 n=5
1. none is defective means r=0
5c0(.2)^0 (.8)^5=0.32768
2. one is defective r=1
5c1(.2)^1(.8)4=5*0.2*.4096
=0.4096
3 r is between 1 and 4 means
r=2,r=3
5c2(.2)^2(0.8)^3 +5c3(.2)^3(0.8)^2
10*.04*.512+10*.008*.64
0.2048+.0512=0.256